OpenStudy (anonymous):

Which of these numbers is a solution to the inequality 4 times a is less than or equal to twelve? 2 4 8 twelve

OpenStudy (stamp):

4 times a = 4*a = 4a is less than or equal to : <= twelve = 12 4a <= 12 4a / 4 <= 12 / 4 a <= 3